# Measuring and Analyzing Intelligence via Contextual Uncertainty in Large Language Models using Information-Theoretic Metrics

### v2

@@ 1 Introduction
Intelligence defies a single, agreed-upon definition Neisser (1979); Sternberg (2004). Yet recent LLMs display behaviours that look intelligent, reaching human-level scores on complex tasks. If the concept itself is elusive Bommasani et al. (2021); Bender et al. (2021), how can we understand systems that seem to embody it?
LLMs give us a unique opening. For any context, they output a full probability distribution over the next token Brown et al. (2020); Bishop and Nasrabadi (2006). These distributions, derived from the model’s internal logits, let us study information processing directly.
We ground our analysis in Shannon’s information theory Shannon (1948). Prior work has used information measures to trace information flow in deep nets Tishby et al. (2000); Shwartz-Ziv and Tishby (2017) and to gauge LLM confidence Kadavath et al. (2022); Jiang et al. (2020). We build on this by noting that intelligent behaviour balances two extremes: rigid determinism, which lacks creativity, and pure randomness, which is noise. A capable system should lock in on clear evidence yet stay flexible when context is thin. We call this capacity Adaptive Predictive Modulation.
@@
$u_{k}:=\frac{h_{k}}{H_{k}},\hskip 17.00024pt0\leq u_{k}\leq 1,$
which places the model on a determinism–randomness scale for each $k$.
By tracing $u_{k}$ against $k$, we obtain the Entropy Decay Curve—a quantitative Cognitive Profile [^1] Perplexity (PPL) relates to conditional entropy via $h=\log_{2}(\text{PPL})$, but PPL is usually reported as a single corpus-level score. The index $u_{k}$ is new: it also uses $H_{k}$, which PPL ignores.. The curve’s starting point, slope, and floor capture a model’s information-handling strategy. We present the first empirical study of these profiles for leading LLMs, showing how they reveal both genuine capability and issues such as data contamination Carlini et al. (2022); Achiam et al. (2023).
@@ 2.1 Predictive Distribution of an LLM
Consider an auto-regressive LLM that, given a context of $k$ tokens, outputs a probability distribution over its full vocabulary. Let
@@
Crucially, $p(Y\mid X)$ refers to the entire probability vector, not just a sampled outcome. The vector is obtained by applying the softmax function to the model’s logits, the standard way to convert raw scores into a valid probability distribution in multi-class generation tasks Goodfellow et al. (2016); Vaswani et al. (2017).
@@ 2.2.2 Marginal Entropy $\boldsymbol{(H_{k})}$
Exact marginal entropy requires averaging over all contexts and is intractable Cover (1999). Instead, we form the average predictive distribution
$\bar{p}_{k}(y):=\frac{1}{N}\sum_{i=1}^{N}p(y\mid X=x_{i}),\hskip 17.00024pty\in\mathcal{Y},$
@@
This value measures the overall diversity of tokens the model finds plausible, given $k$-token contexts, and serves as an upper bound for $h_{k}$.
@@ 3.1 Models
  • DeepSeek-R1 8.19B — 8.19-billion parameters in the Qwen3 line Yang et al. (2025);
  • Qwen 2.5 7.62B — 7.62-billion parameters in the Qwen2 family Yang et al. (2024).
@@ 3.3 Implementation Details and Procedure
      Compute the conditional entropy $h_{x_{i}}$; store $p(Y\mid X=x_{i})$.
@@
4 Results
We now quantify the predictive uncertainty exhibited by the three language models across multiple context lengths and corpora. The analysis centres on the conditional entropy $h_{k}$, the marginal entropy $H_{k}$, and their ratio $u_{k}$.
@@ 4.1 Empirical Data and Initial Observations
Tables 1, 2, and 3 report $h_{k}$, $H_{k}$, and the derived uncertainty index $u_{k}$ for all models on Ulysses, Alice’s Adventures in Wonderland, and Kant’s Critique of Judgement, respectively. The primary metric $u_{k}$ expresses residual prediction uncertainty normalised by the model’s potential output diversity at each context length $k$.
@@
Monotonic Decline in Uncertainty
For every model and corpus, the uncertainty index $u_{k}$ decreases as the context length $k$ grows. Models are most uncertain at $k=3$; confidence rises at $k=9$ and continues to improve through $k=30$ and beyond. This uniform trend confirms that additional context restricts the set of plausible next tokens.
@@
Although $u_{k}$ falls steadily, its decline slows once the context becomes long. The drop from $k=3$ to $k=30$ is much larger than the drop from $k=300$ to $k=600$. Thus most predictive benefit comes from early context, while very long windows add only modest gains.
@@ 4.2 Model-Specific Cognitive Profiles
Figures 1, 2, and 3 plot the Entropy Decay Curves for each model–corpus pair. The $x$-axis is logarithmic to highlight behaviour at short context lengths. These curves reveal how each architecture leverages context and where their strategies diverge.
Figure 1: Entropy Decay Curves ($u_{k}$ vs. $k$) for all three models on the “Alice’s Adventures in Wonderland” corpus. The horizontal axis is shown on a logarithmic scale.
Figure 2: Entropy Decay Curves ($u_{k}$ vs. $k$) for all three models on the “Ulysses” corpus. The horizontal axis is shown on a logarithmic scale.
Figure 3: Entropy Decay Curves ($u_{k}$ vs. $k$) for all three models on the “Kant’s Critique of Judgement” corpus. The horizontal axis is shown on a logarithmic scale.
  • Large-Scale Model (Llama 3.3). At very short contexts ($k=3$) Llama 3.3 shows the highest initial uncertainty, signalling a wide search over possible continuations. Its curve then drops steeply, reaching the lowest uncertainty once the context exceeds about $k=90$. This rapid shift from broad exploration to sharp focus appears to be a hallmark of its large parameter count.
  • Smaller Models (DeepSeek-R1 and Qwen 2.5). The two smaller models display a gentler decline. On challenging texts such as Ulysses and Kant’s Critique of Judgement, their curves remain relatively flat and end above $u_{k}=0.35$. This suggests a limited ability to exploit long-range context compared with the larger model.
@@
The uncertainty index $u_{k}$ varies strongly with the corpus. For all models, Alice’s Adventures in Wonderland yields the lowest $u_{k}$ values, indicating that this text is highly predictable. In contrast, Ulysses and Kant’s Critique produce higher $u_{k}$ values, reflecting their greater difficulty for next-token prediction. These differences show that the Entropy Decay Curve (EDC) captures how model capability interacts with linguistic complexity.
@@ 4.3 Detecting Memorisation and Data Contamination
The EDC also flags potential memorisation. When Llama 3.3 is run on the Alice corpus, its $u_{k}$ values fall almost to zero at long contexts, an extremity unlikely to arise from genuine generalisation. This strongly suggests that large portions of Alice appeared in its pre-training data. By contrast, on the Kant corpus the same model keeps noticeably higher $u_{k}$ values, consistent with real predictive processing. Thus the EDC can (i) audit test sets for contamination and (ii) separate models that truly generalise from those that rely mainly on memorised content.
@@ 4.4 The Information Gain Span (IGS) as a Summary Metric
Table 4: Information Gain Span (IGS) Index for Each Model and Corpus. A higher IGS score indicates a better balance between high initial uncertainty and low final uncertainty.
Model Corpus
Alice Ulysses Kant
Llama 3.3 70.6B 0.6764 0.6523 0.5836
DeepSeek-R1 8.19B 0.4339 0.3161 0.3492
Qwen2.5 7.62B 0.5176 0.3225 0.3544
The IGS scores quantitatively confirm our qualitative observations. The Llama 3.3 model achieves the highest IGS score on the Alice and Ulysses corpora, reflecting its superior ability to transition from divergent to convergent processing. All models score lower on the more complex texts, quantitatively demonstrating the increased difficulty.
@@ 5 Discussion
The Entropy Decay Curves (EDCs) do more than rank models; they provide a direct, quantitative view of how large language models process information. Because each model–corpus pair yields a distinct curve, the results illuminate properties of both the models and the texts. We highlight two main implications.
5.1 Predictability Profiles of Model–Text Interaction
An EDC serves as a predictability profile: it shows how quickly a model’s uncertainty falls as it reads a text. For every model tested, Ulysses and Kant’s Critique of Judgement produce higher curves than Alice’s Adventures in Wonderland. The first two works therefore pose a longer-lasting predictive challenge.
Unlike traditional readability indices Kincaid et al. (1975), this profile is data-driven. The model itself reveals a text’s effective predictability through its evolving output distribution, without relying on hand-crafted complexity rules.
@@ 5.2 Entropy Collapse as an Anomaly Signal
The same method flags unusual behaviour. On the Alice corpus, Llama 3.3 pushes the uncertainty index $u_{k}$ almost to zero at long contexts—a phenomenon we call entropy collapse. Such extreme certainty is unlikely without exposure to the exact text and is therefore strong evidence of memorisation from training-data contamination Carlini et al. (2022); Achiam et al. (2023). Because Alice’s Adventures in Wonderland is a public-domain classic, its presence in pre-training data is plausible.
@@ 6 Theoretical Properties of the Entropy Metrics
This section provides a theoretical basis for the empirically observed patterns of $h_{k}$ and $H_{k}$. We move from sample averages to expectations taken with respect to the model’s full probability distribution $p$.
@@ 6.1 From Sample Estimates to Expected Values
Definition 1 (Theoretical Entropy Metrics).
The theoretical average conditional entropy is
$h_{k}^{\text{theory}}\;:=\;\mathbb{E}_{X_{1:k}\sim p}\!\bigl[H\!\bigl(p(Y_{k+1}\mid X_{1:k})\bigr)\bigr].$
The theoretical entropy of the average distribution is
$H_{k}^{\text{theory}}\;:=\;H\!\Bigl(\,\mathbb{E}_{X_{1:k}\sim p}\!\bigl[p(Y_{k+1}\mid X_{1:k})\bigr]\Bigr).$
Because the inner expectation is the marginal distribution of $Y_{k+1}$, we have
$H_{k}^{\text{theory}}\;=\;H\!\bigl(p(Y_{k+1})\bigr).$
The empirical $h_{k}$ and $H_{k}$ are Monte Carlo estimates of these theoretical quantities. For the rest of this section we drop the superscript “theory” for clarity.
@@ 6.2 Monotonicity of the Average Conditional Entropy ($h_{k}$)
Empirical evidence shows that the average conditional entropy $h_{k}$ does not increase as the context length $k$ grows. A formal proof is subtler, because language is non-stationary and a direct appeal to the inequality $H(Y\mid X,A)\leq H(Y\mid X)$ would compare different prediction targets ($Y_{k+1}$ versus $Y_{k+2}$). To bridge this gap, we first state a sufficient, non-trivial condition that well-trained language models are expected to satisfy. It captures the idea that a model’s confidence, given a fixed context, should not improve when predicting events further in the future.
Proposition 1 (Monotonicity of $h_{k}$ under Predictive Attenuation).
Let $p$ be an autoregressive model. Assume the Predictive Attenuation Condition: for every $k\geq 0$,
$\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+2}\mid X_{1:k})\bigr]\;\leq\;\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+1}\mid X_{1:k})\bigr].$ (2)
Then $h_{k+1}\leq h_{k}$.
@@
Step 1. For any fixed context $X_{1:k+1}$,
$H(Y_{k+2}\mid X_{1:k},X_{k+1})\;\leq\;H(Y_{k+2}\mid X_{1:k}).$
Taking expectations over $X_{1:k+1}$ yields
$h_{k+1}\;\leq\;\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+2}\mid X_{1:k})\bigr].$ (3)
Step 2. Applying (2),
$\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+2}\mid X_{1:k})\bigr]\;\leq\;\mathbb{E}_{X_{1:k}}\!\bigl[H(Y_{k+1}\mid X_{1:k})\bigr]\;=\;h_{k}.$ (4)
Combining (3) and (4) gives $h_{k+1}\leq h_{k}$. ∎
Why the Condition Is Reasonable.
Equation (2) states that, on average, a fixed context tells the model less about a token two steps ahead than about the next token. Because the unobserved variable $Y_{k+1}$ introduces additional uncertainty, this behaviour is expected in well-trained language models.

### v3

@@ Abstract
keywords:
Neural Networks, Large Language Models, Information Theory, Conditional Entropy, Uncertainty Quantification, Memorisation
@@ 1 Introduction
Intelligence defies a single, agreed-upon definition [neisser1979concept, sternberg2004culture]. Yet recent LLMs display behaviours that look intelligent, reaching human-level scores on complex tasks. If the concept itself is elusive [bommasani2021opportunities, bender2021dangers], how can we understand systems that seem to embody it?
LLMs give us a unique opening. For any context, they output a full probability distribution over the next token [brown2020language, bishop2006pattern]. These distributions, derived from the model’s internal logits, let us study information processing directly.
We ground our analysis in Shannon’s information theory [shannon1948]. Prior work has used information measures to trace information flow in deep nets [tishby2000information, shwartz2017opening] and to gauge LLM confidence [kadavath2022language, jiang2020can]. We build on this by noting that intelligent behaviour balances two extremes: rigid determinism, which lacks creativity, and pure randomness, which is noise. A capable system should lock in on clear evidence yet stay flexible when context is thin. We call this capacity Adaptive Predictive Modulation.
@@
$u_{k}:=\frac{h_{k}}{H_{k}},\qquad 0\leq u_{k}\leq 1,$
which places the model on a determinism—randomness scale for each $k$.
By tracing $u_{k}$ against $k$, we obtain the Entropy Decay Curve—a quantitative Cognitive Profile [^1] Perplexity is typically derived from cross-entropy on ground-truth next tokens, whereas $h_{k}$ here is the predictive entropy of the model’s full next-token distribution. Our index $u_{k}$ additionally normalises by $H_{k}$, which PPL does not capture.. The curve’s starting point, slope, and floor capture a model’s information-handling strategy. We present an empirical study of these profiles for several leading LLMs, showing how they reveal both genuine capability and issues such as data contamination [carlini2022quantifying, achiam2023gpt].
@@ 2.1 Predictive Distribution of an LLM
Consider an autoregressive LLM that, given a context of $k$ tokens, outputs a probability distribution over its full vocabulary. Let
@@
Crucially, $p(Y\mid X)$ refers to the entire probability vector, not just a sampled outcome. The vector is obtained by applying the softmax function to the model’s logits, the standard way to convert raw scores into a valid probability distribution in multi-class generation tasks [goodfellow2016deep, vaswani2017attention].
@@ 2.2.2 Marginal Entropy $\boldsymbol{(H_{k})}$
Exact marginal entropy requires averaging over all contexts and is intractable [cover1999elements]. Instead, we form the average predictive distribution
$\bar{p}_{k}(y):=\frac{1}{N}\sum_{i=1}^{N}p(y\mid x_{i}),\qquad y\in\mathcal{Y},$
@@
By concavity of entropy,
$H_{k}=H(\bar{p}_{k})\;\geq\;\frac{1}{N}\sum_{i=1}^{N}H\!\bigl(p(\cdot\mid x_{i})\bigr)\;=\;\frac{1}{N}\sum_{i=1}^{N}h_{x_{i}}\;=\;h_{k},$
so $H_{k}$ serves as an upper bound for $h_{k}$.
@@ 3.1 Models
  • DeepSeek-R1 8.19B — 8.19-billion parameters in the Qwen3 line [qwen3];
  • Qwen 2.5 7.62B — 7.62-billion parameters in the Qwen2 family [qwen2.5].
@@ 3.3 Implementation Details and Procedure
      Compute the conditional entropy $h_{x_{i}}$; accumulate $p(Y\mid X=x_{i})$ into a running average to form $\bar{p}_{k}$.
@@
3.4 Positional Robustness Check (Positional Homogeneity)
A potential concern in sliding-window entropy estimation is that averages at different configurations may implicitly mix different target positions in the source text, and thus reflect positional non-stationarity rather than a pure context-length effect. To empirically probe this issue, we ran blockwise positional robustness checks on the Alice’s Adventures in Wonderland corpus using Llama 3.3 at two context lengths: a short window ($k=3$) and a long window ($k=600$).
We select three disjoint target blocks indexed by $j_{0}\in\{600,1200,1800\}$. For each block and each $k\in\{3,600\}$ we form
$X^{(k)}_{j}=(t_{j-k},t_{j-k+1},\dots,t_{j-1}),\qquad Y_{j}=t_{j},\qquad j\in\{j_{0},\dots,j_{0}+999\},$
compute the conditional entropy $H(Y_{j}\mid X^{(k)}_{j})$ and the corresponding predictive distribution $p(\cdot\mid X^{(k)}_{j})$ for $N=1000$ targets, then aggregate to obtain $h_{k}$, $H_{k}$, and $u_{k}=h_{k}/H_{k}$.
Table 1: Positional robustness on Alice’s Adventures in Wonderland for Llama 3.3 at two fixed context lengths ($k=3$ and $k=600$). Each row averages over $N=1000$ consecutive targets in the indicated range. Reported values are the mean conditional entropy $h_{k}$, the entropy of the average predictive distribution $H_{k}$, and the uncertainty index $u_{k}=h_{k}/H_{k}$.
Target range $j$ $k=3$ $k=600$
$\boldsymbol{h_{3}}$ (bits) $\boldsymbol{H_{3}}$ (bits) $\boldsymbol{u_{3}}$ $\boldsymbol{h_{600}}$ (bits) $\boldsymbol{H_{600}}$ (bits) $\boldsymbol{u_{600}}$
600–1599 11.5641 13.3657 0.8652 0.1161 7.7403 0.0150
1200–2199 11.6408 13.4568 0.8650 0.1194 7.7721 0.0154
1800–2799 11.7438 13.4985 0.8700 0.1323 7.7415 0.0171
Range (max–min) 0.1797 0.1328 0.0050 0.0162 0.0318 0.0021
Across these three within-text blocks, the uncertainty index remains stable at both short and long context lengths as in Table 1. For $k=3$, $u_{3}$ varies only from $0.8650$ to $0.8700$ (absolute range $0.0050$, i.e. $\approx 0.58\%$ relative to the minimum). For $k=600$, $u_{600}$ varies from $0.0150$ to $0.0171$ (absolute range $0.0021$). Overall, these results suggest that within the tested portion of the corpus, positional effects are small relative to the magnitude of the entropy-based quantities, supporting the use of block-averaged estimates for the main experiments. These positional checks are auxiliary control experiments and therefore use a longer contiguous segment of the corpus than the 1600-token segment used in the main EDC measurements.
@@ 3.5 Empirical Data and Initial Observations
Tables 2, 3, and 4 report $h_{k}$, $H_{k}$, and the derived uncertainty index $u_{k}$ for all models on Ulysses, Alice’s Adventures in Wonderland, and Kant’s Critique of Judgement, respectively. The primary metric $u_{k}$ expresses residual prediction uncertainty normalised by the model’s potential output diversity at each context length $k$.
@@
Overall Decline in Uncertainty
Across all model-corpus pairs, the uncertainty index $u_{k}$ generally decreases as context length $k$ grows, with the largest reductions occurring between short windows (e.g., $k=3$ to $k=30$). At the longest windows ($k=300$ to $k=600$), some runs exhibit small non-monotonic fluctuations, consistent with finite-sample variance in the Monte Carlo estimates.
@@
Although $u_{k}$ falls steadily, its decline slows once the context becomes long. The change from $k=300$ to $k=600$ is small and can be non-monotonic, consistent with finite-sample variance in the Monte Carlo estimates. Thus most predictive benefit comes from early context, while very long windows add only modest gains.
@@ 3.6 Model-Specific Cognitive Profiles
Figures 13 plot the Entropy Decay Curves for each model-corpus pair. The $x$-axis is logarithmic to emphasise behaviour at short context lengths.
Figure 1: Entropy Decay Curves ($u_{k}$ vs. $k$) on Alice’s Adventures in Wonderland.
Figure 2: Entropy Decay Curves ($u_{k}$ vs. $k$) on Ulysses.
Figure 3: Entropy Decay Curves ($u_{k}$ vs. $k$) on Kant’s Critique of Judgement.
  • Large-scale model (Llama 3.3). At very short contexts ($k=3$), Llama 3.3 exhibits the highest uncertainty across the tested corpora, indicating a broad set of plausible continuations. As $k$ increases, its uncertainty decays much more rapidly than the smaller models, reaching the lowest long-context levels (notably on Alice and Ulysses). This sharp transition is consistent with stronger context utilisation at larger scale.
  • Smaller models (DeepSeek-R1 and Qwen 2.5). Both smaller models show a gentler decline in $u_{k}$. On challenging texts such as Ulysses and Kant’s Critique of Judgement, their curves remain comparatively high even at long contexts, suggesting more limited long-range disambiguation.
@@
The uncertainty index $u_{k}$ depends strongly on the corpus. Across all three models, Alice’s Adventures in Wonderland exhibits substantially lower uncertainty at moderate-to-long context lengths (notably $k\geq 30$). This implies that once sufficient conditioning context is available, the model’s next-token distribution becomes sharply concentrated. By contrast, Ulysses and Kant’s Critique of Judgement maintain higher $u_{k}$ over the same range, indicating that considerable residual uncertainty persists even with long contexts. In this sense, the Entropy Decay Curve (EDC) functions as an empirical predictability profile: it quantifies how effectively a model achieves contextual disambiguation across texts of differing effective predictive difficulty.
@@ 3.7 Detecting Memorisation and Data Contamination
The EDC also flags potential memorisation. When Llama 3.3 is run on the Alice corpus, its $u_{k}$ values fall almost to zero at long contexts, an extremity unlikely to arise from genuine generalisation. This is consistent with the hypothesis that large portions of Alice appeared in its pre-training data. By contrast, on the Kant corpus the same model keeps noticeably higher $u_{k}$ values, consistent with real predictive processing. Thus the EDC can (i) audit test sets for contamination and (ii) separate models that truly generalise from those that rely mainly on memorised content.
@@ 3.8 The Information Gain Span (IGS) as a Summary Metric
Table 5: Information Gain Span (IGS) computed as $\mathrm{IGS}=u_{3}(1-u_{600})$. A higher IGS score indicates a better balance between high initial uncertainty and low final uncertainty.
Model Alice Ulysses Kant
Llama 3.3 70.6B 0.8539 0.8054 0.6942
DeepSeek-R1 8.19B 0.4707 0.3470 0.3715
Qwen2.5 7.62B 0.5745 0.3548 0.3773
The IGS scores quantitatively confirm our qualitative observations. The Llama 3.3 model achieves the highest IGS score on the Alice and Ulysses corpora, reflecting its superior ability to transition from divergent to convergent processing. IGS is highest on Alice for all models, while Ulysses and Kant yield lower and broadly comparable scores, with their ordering varying by model.
@@ 4 Discussion
The Entropy Decay Curves (EDCs) do more than rank models; they provide a direct, quantitative view of how large language models process information. Because each model-corpus pair yields a distinct curve, the results illuminate properties of both the models and the texts. We highlight two main implications.
4.1 Predictability Profiles of Model-Text Interaction
An EDC serves as a predictability profile: it shows how quickly a model’s uncertainty falls as it reads a text. For every model tested, Ulysses and Kant’s Critique of Judgement yield higher uncertainty than Alice for moderate-to-long contexts (especially $k\geq 30$), and therefore remain predictive challenges over longer spans. The first two works therefore pose a longer-lasting predictive challenge.
Unlike traditional readability indices [kincaid1975derivation], this profile is data-driven. The model itself reveals a text’s effective predictability through its evolving output distribution, without relying on hand-crafted complexity rules.
@@ 4.2 Entropy Collapse as an Anomaly Signal
The same method flags unusual behaviour. On the Alice corpus, Llama 3.3 pushes the uncertainty index $u_{k}$ almost to zero at long contexts—a phenomenon we call entropy collapse. Such extreme certainty is unlikely without exposure to the exact text and is therefore strong evidence of memorisation from training-data contamination [carlini2022quantifying, achiam2023gpt]. Because Alice’s Adventures in Wonderland is a public-domain classic, its presence in pre-training data is plausible.
@@ 5 Theoretical Properties of the Entropy Metrics
This section provides a theoretical basis for the empirically observed patterns of $h_{k}$ and $H_{k}$. We relate the finite-sample estimators of Section 2 to corresponding population quantities under a local sampling distribution over contexts.
@@ 5.1 From Sample Estimates to Expected Values
We now define population (theoretical) counterparts of the empirical quantities $h_{k}$ and $H_{k}$ by replacing the finite-window average by an expectation under the same sliding-window sampling rule.
Fix a context length $k$ and consider the evaluated token sequence $(T_{1},T_{2},\dots)$. Select a window-start index $I$ uniformly at random from the admissible starts used by the estimator (e.g. $I\in\{1,\dots,N\}$ in our experiments). Define the random context and target
$X^{(k)}:=(T_{I},\dots,T_{I+k-1}),\qquad Y^{(k)}:=T_{I+k}.$
Given a realised context $X^{(k)}=x$, the model outputs next-token probabilities $p(y\mid x)$ for $y\in\mathcal{Y}$. Using the same predictive-entropy definition as in Section 2, define
$H(Y^{(k)}\mid X^{(k)}=x)\;:=\;-\sum_{y\in\mathcal{Y}}p(y\mid x)\,\log_{2}p(y\mid x).$
Theoretical conditional entropy.
We define the theoretical average conditional entropy at length $k$ as
$h_{k}\;:=\;\mathbb{E}\!\left[\,H\!\left(Y^{(k)}\mid X^{(k)}\right)\right]\;=\;\mathbb{E}\!\left[\,-\sum_{y\in\mathcal{Y}}p\!\left(y\mid X^{(k)}\right)\log_{2}p\!\left(y\mid X^{(k)}\right)\right],$
where the expectation is over the random window start $I$ (equivalently, over the induced random window $(X^{(k)},Y^{(k)})$).
Theoretical marginal entropy.
Define the averaged predictive distribution
$\bar{p}_{k}(y)\;:=\;\mathbb{E}\!\left[p\!\left(y\mid X^{(k)}\right)\right],\qquad y\in\mathcal{Y},$
and its entropy
$H_{k}\;:=\;-\sum_{y\in\mathcal{Y}}\bar{p}_{k}(y)\,\log_{2}\bar{p}_{k}(y).$
Connection to the empirical estimators.
The empirical quantities of Section 2 are Monte Carlo estimates of these expectations under the uniform sliding-window sampling rule: the sample mean $\frac{1}{N}\sum_{i}$ approximates $\mathbb{E}[\cdot]$, and the empirical average distribution $\frac{1}{N}\sum_{i}p(y\mid x_{i})$ approximates $\bar{p}_{k}(y)$.
@@ 5.2 Monotonicity of the Average Conditional Entropy ($h_{k}$)
Empirically, $h_{k}$ decreases as $k$ grows. In our setting, $h_{k}$ and $h_{k+1}$ correspond to different targets ($T_{I+k}$ versus $T_{I+k+1}$), so we use a local shift-invariance assumption to compare them under the same window-averaging rule.
Assumption 1 (Positional homogeneity / local shift-invariance).
Under the sliding-window sampling rule, the distribution of local blocks is (approximately) invariant to a one-step shift: replacing every sampled block $(T_{I},\dots,T_{I+k})$ by $(T_{I-1},\dots,T_{I+k-1})$ does not materially change its distribution within the evaluated segment.
Proposition 1 (Monotonicity of $h_{k}$ under positional homogeneity).
Assume Assumption 1. With
$h_{k}:=\mathbb{E}\!\left[\,H\!\left(T_{I+k}\mid T_{I},\dots,T_{I+k-1}\right)\right],$
we have $h_{k+1}\leq h_{k}$ (and, under approximate homogeneity, $h_{k+1}\lesssim h_{k}$).
@@
By definition,
$h_{k+1}=\mathbb{E}\!\left[\,H\!\left(T_{I+k+1}\mid T_{I},\dots,T_{I+k}\right)\right].$
By Assumption 1, shifting indices by one gives
$h_{k+1}\;\approx\;\mathbb{E}\!\left[\,H\!\left(T_{I+k}\mid T_{I-1},\dots,T_{I+k-1}\right)\right].$
Conditioning reduces entropy pointwise, hence
$H\!\left(T_{I+k}\mid T_{I-1},\dots,T_{I+k-1}\right)\;\leq\;H\!\left(T_{I+k}\mid T_{I},\dots,T_{I+k-1}\right).$
Taking expectations yields $h_{k+1}\lesssim h_{k}$, and under exact shift-invariance the inequality is $h_{k+1}\leq h_{k}$. ∎
Connection to our estimator.
For each realised window start $I=i$, the term $H(T_{i+k}\mid T_{i},\dots,T_{i+k-1})$ is exactly the predictive entropy computed from the model’s next-token distribution $p(y\mid T_{i},\dots,T_{i+k-1})$ as in Section 2. Averaging over $I$ matches the sliding-window averaging used to compute the empirical $h_{k}$.